how who u go about sloving 40 centimeters per second to miles per hour.
and 1000 cubic centimeters to cubic feet

Answer by Alan3354(69443)

how who u go about sloving 40 centimeters per second to miles per hour.
and 1000 cubic centimeters to cubic feet
---------
It's not solving, it's converting.
1 inch = 2.54 cm exactly
--------
40 cm/sec = 40*60*60 cm/hour = 144000cm/hour
= 144000/2.54 inch/hour
= (144000/2.54)/(12*5280) miles/hr
=~ 0.89477 mph
----------------
1000 cc/(2.54*2.54*2.54) = cubic inches
= (1000/2.54^3)/12^3 cubic ft
=~ 0.0353 ft^3